Position of a particle varies with time x=t^2-2x+1 then the magnitude of displacement of the particle at t=1s to t=2s

Grade:11

1 Answers

Venkat
5 years ago
At t = 1 , x =1-2x+1 → 3x = 2 → x = 3/2
at t = 2s , 3x =2^1+1  → 3x = 5 x → x = 5/3
displacement = 5/3 – 3/2 = 4/6
typo error . please ignore above 
At t = 1 , x =1-2x+1 → 3x = 2 → x = 3/2
at t = 2s , 3x =2^1+1  → 3x = 5 x → x = 5/3
displacement = 5/3 – 3/2 = 1/6